# Learning Equilibria from Data: Provably Efficient Multi-Agent Imitation Learning

### v1

@@ Appendix F Analysis of BR Oracle Algorithm
where $\mu^{\star}_{k}\in\mathrm{br}(\nu_{k})$ and $\nu^{\star}_{k}\in\mathrm{br}(\mu_{k})$. At this point, dividing by $K$ squaring and applying the performance difference Lemma L.1 that allows to decompose the global regret into a weighted some of regrets at each state, we obtain
@@
where the second last step used the Cauchy-Schwartz inequality and the last step used the Jensen’s inequality and the concavity of the square root. Analogous steps give
@@
Therefore, thanks to an application of the Azuma-Hoeffding inequality it holds that with probability $1-\delta$
@@
where for the first term, we recognized that the policies updates in Algorithm 1 can be seen as mirror descent updates (see Lemma F.1) and we used the standard regret bound for online mirror descent (see for example Orabona [2023]) instantiated with the following Bregman divergence $\sum_{s\in{\mathcal{S}}}D_{KL}(\mu(\cdot|s),\mu^{\prime}(\cdot|s))$ and with learning rate $\eta=\frac{2\left|{{\mathcal{S}}}\right|\log\left|{{\mathcal{A}}_{\max}}\right|}{K}$ as done in Algorithm 1. All in all, we obtain via a union bound that with probability at least $1-4\delta$
@@
The next results shows that the policies updates used in Algorithm 1 and Algorithm 2 are mirror descent updates for an appropriately chosen Bregman divergence. To this end for any $p,d\in\Delta_{\mathcal{A}}$ we define the KL divergence as $KL(p,q)=\sum_{a\in\mathcal{A}}p(a)\log(p(a)/q(a))$ with the convention that $KL(p,q)=0$ if there exists an action $a$ such that $p(a)=0$ and $q(a)>0$.
Lemma F.1.
The updates used in Algorithm 1 and Algorithm 2, that is
@@
The Bregamn divergence chosen is induced by the function $\psi(\mu)=\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\mu(a|s)\log\mu(a|s)$ sum of the negative entropy over the state space. Notice that the gradient norm tends to infinite as $\mu$ approaches the border of the policy space (i.e. some entries $\mu(a|s)$ tends to zero), that is $\lim_{\mu\rightarrow\delta\Pi}\left\|{\nabla\psi(\mu)}\right\|$. This means that the first order optimality condition implies that the derivative $F(\mu)=\left\langle{g^{\mu}_{k}},{\mu}\right\rangle+\frac{1}{\eta}\sum_{s\in{\mathcal{S}}}KL(\mu(\cdot|s),\mu_{k}(\cdot|s))$ equals zero at the minimizing policy which is $\mu_{k+1}$ by definition. Therefore, in the following we use this fact to derive the exponential weight updates used in Algorithm 1 and 2.
@@
for some $c\in\mathbb{R}$ which ensures normalization of $\mu_{k+1}$. Therfore, inverting the last expression, we have that
@@
Choosing $c\in\mathbb{R}$ to ensure that for all $s\in\mathcal{S}$ it holds that $\sum_{a\in\mathcal{A}}\mu_{k+1}(a|s)=1$ concludes the proof. ∎
Appendix G Analysis of Algorithm 2
This section presents the analysis of Algorithm 2 which provides a sample complexity guarantee without requiring neither concentrability or best response oracle. See 4.2
@@ Appendix G Analysis of Algorithm 2
The proof follows similar to the one of Theorem 4.1 with the addition of an RL inner loop. In a first step, we first derive the same decomposition as in (5). Again, dividing by $K$ squaring, applying the performance difference, the Cauchy-Schwartz inequality leads to (6) and using Jensen’s inequality and the concavity of the square root we get
@@
where $\mu^{\star}_{k}\in\mathrm{br}(\nu_{k}).$ At this point, as we do not assume to have access to a Best Response oracle, let us introduce the sequence $\left\{{z_{k}}\right\}^{K}_{k=1}$ and $\left\{{y_{k}}\right\}^{K}_{k=1}$ produced by UCB-VI in the inner loop of Algorithm 2. Since the stochastic reward used in the inner loop is unbiased and almost surely bounded by $2$ by Lemma H.7, Lemma H.6 run for a number of inner iterations $T=\mathcal{O}\left({\frac{\left|{{\mathcal{S}}}\right|^{3}\left|{{\mathcal{A}}_{\max}}\right|\log(1/\delta)}{(1-\gamma)^{4}\varepsilon^{2}_{\mathrm{opt}}}}\right)$ ensures that with probability $1-3\delta$
@@
Note that now we can sample $S_{k}^{\mu}\sim d^{\mu_{k},y_{k}}$ and $S_{k}^{\nu}\sim d^{z_{k},\nu_{k}}$. Therefore, we can again add and subtract the terms $\left\|{\mu_{E}(\cdot|S^{\mu}_{k})-\mu_{k}(\cdot|S^{\mu}_{k})}\right\|^{2}$ in (G) and $\left\|{\nu_{E}(\cdot|S^{\nu}_{k})-\nu_{k}(\cdot|S^{\nu}_{k})}\right\|^{2}$ in (9) we obtain that
@@
We have that (12) can be bounded analogously as in Eq. Martingale with $S^{\mu}_{k}\sim d^{\mu_{k},y_{k}}.$
Again, as done in the proof of Theorem 4.1, we recognize that the policies updates performed by Algorithm 2 are instances of online mirror descent ( see Lemma F.1 ). Therefore, we can bound the regret term as follows
@@
Appendix H Analysis for the RL inner loop
@@
and we prove that under the above conditions pointwise optimism hold. This point is made precise in the next Lemma
Lemma H.2.
@@
Using the point wise optimism in Lemma H.2 and the decomposition in Lemma H.1 we have the following decomposition on the regret of UCBVI
@@
$=\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{Q_{k+1}-r+\gamma PV_{t}}\right\rangle+\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{Q_{t}-Q_{k+1}}\right\rangle$
$\leq\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{\widehat{r}_{t}+\gamma\widehat{P}_{t}V_{t}+b_{t}-r+\gamma PV_{t}}\right\rangle+\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{Q_{t}-Q_{k+1}}\right\rangle$
$\leq\frac{2}{K}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{b_{t}}\right\rangle+\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{Q_{t}-Q_{t+1}}\right\rangle$
@@
$\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{Q_{t}-Q_{k+1}}\right\rangle$ $\leq\frac{1}{T}\sum^{T}_{t=1}\sum_{s,a}Q_{t}(s,a)-Q_{t+1}(s,a)$
@@
H.1 Showing validity of the bonuses
@@
For all $t\in[T]$ simultaneously, we have the following upper bound high probability upper bound
@@
where the last inequality follows with probabilitry $1-\delta$ from an application of the Azuma Hoeffding inequality making special care of the fact that the total number of visits $N_{t}(s,a)$ is not an independent random variable with respect to the random variables of which we are computing the mean, that is $\left\{{\mathds{1}_{\left\{{S_{\tau},A_{\tau}=s,a}\right\}}}\right\}^{T}_{t=1}$. For this reason we pay the factor $\log(N_{t}(s,a)(N_{t}(s,a)+1))$ in the upper bound. We refer the reader to [Lattimore and Szepesvári, 2020, Exercise 7.1 ] for details. Since, we have also Therefore, by triangular inequality and a union bound over the state space.
@@
H.2 Bound the bonus sum
Lemma H.5.
@@ H.2 Bound the bonus sum
H.3 Final UCBVI bound
Lemma H.6.
Let us consider UCBVI (Algorithm 3) in an environment with a stochastic unbiased reward almost surely bounded by $2$ run for $T$ iteration with a sequence of valid bonuses $\left\{{b_{t}}\right\}^{T}_{t=1}$ specified in the statement of Lemma H.4, then it holds that with probability at least $1-3\delta$
@@ H.3 Final UCBVI bound
The proof follows trivially from the combination of Lemma H.5 and Lemma H.3 and a union bound over the event that the bonus are valid and the event under which the bound in Lemma H.5 holds .
H.4 Properties of the reward estimate
Lemma H.7.
For any policy $\pi\in\Pi$ and expert policy $\pi_{E}\in\Pi$ consider a particular state $s\in{\mathcal{S}}$ and sampling $A_{E}\sim\pi_{E}(\cdot|s)$, $A^{\prime}_{E}\sim\pi_{E}(\cdot|s)$ Then, the following facts hold true
@@ H.4 Properties of the reward estimate
$\left\|{x_{E}-\pi}\right\|^{2}$ $=\left\|{x_{E}(\cdot|s)}\right\|^{2}-2\left\langle{x_{E}(\cdot|s)},{x^{k}(\cdot|s)}\right\rangle+\left\|{\pi(\cdot|s)}\right\|^{2}$
$=\sum_{a}x_{E}(a\mid s)^{2}+\sum_{a}\pi(a\mid s)^{2}-2\sum_{a}x_{E}(a\mid s)\pi(a\mid s)$
$=\sum_{a}x_{E}(a\mid s)^{2}+\sum_{a}\pi(a\mid s)^{2}-2\mathbb{E}_{A\sim x_{E}(\cdot\mid s)}[\pi(a\mid s)].$
@@
$x_{E}^{2}(a\mid s)=\mathbb{P}(A_{E}=a)^{2}=\mathbb{P}(A_{E}=a)\mathbb{P}(A^{\prime}_{E}=a)=\mathbb{P}(A_{E}=A^{\prime}_{E})=\mathbb{E}[\mathds{1}_{\{A_{E}=A^{\prime}_{E}\}}],$
where $A_{E},A^{\prime}_{E}$ are independent samples from $x_{E}(\cdot\mid s).$ Therefore, we can conclude that
@@
is an unbiased estimator of $\left\|{x_{E}-\pi}\right\|^{2}$. The second statement is easy to show
@@
Appendix I Extension to $n$-player general-sum Games
In this section, we sketch the analysis for the $n$-player general sum extension. The goal of this section is to show that the algorithm design is decentralized, meaning that it can avoid the curse of multi-agents in $n$-player general-sum Games as stated in Remark 4.1. The idea is that the introduced algorithms keep the other players fixed, in the RL inner-loop and the BR oracle calls respectively. This results in a decentralized execution. This section starts with the introduction of $n$-player general-sum Games and all the necessary notations. Then, we show how the objective varies slightly from the one in Zero-Sum Games. Last, we give the proof sketch for the $n$- player general-sum case.
@@ Appendix I Extension to $n$-player general-sum Games
$V_{i}^{\boldsymbol{\pi}}(s):=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}r_{i}(s,\textbf{a})\mid s_{0}=s\right]$
$Q_{i}^{\boldsymbol{\pi}}(s,\textbf{a}):=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}r_{i}(s,\textbf{a})\mid s_{0}=s,a_{0}=\textbf{a}\right].$
All other expressions are defined as in Section 2 with the extension that the joint actions are now given by a and one fixes the polices of all players expect player $i$, i.e. $\pi_{-i}$, for the induced Games.
@@
$\leq\sum_{i=1}^{n}\underbrace{V^{\pi^{\star}_{i},{\pi}_{-i}}_{i}(s_{0})-V^{\pi^{\star}_{i},{\pi}_{E_{-i}}}_{i}(s_{0})}_{\mathrm{Exploit-Gap}_{i}}+\underbrace{V^{\pi_{E_{i}},{\pi}_{E_{-i}}}_{i}(s_{0})-V^{\pi_{i},{\pi}_{-i}}_{i}(s_{0})}_{\mathrm{Value-Gap}}.$
The Exploit-Gap is similar to the objective analyzed in the zero-sum case with the difference that now that the policies of the other players are varying in $n-1$ cases. The Value-Gap is new and does not appear in the zero-sum case as one can again use the structure of the reward in that case. However, the latter is easy to analyze as it can be seen as a single-agent MDP with the joint policy $\boldsymbol{\pi}$ and the joint expert $\boldsymbol{\pi}_{E}$ respectively.

### v2

@@ Appendix E Analysis of BR Oracle Algorithm
where $\mu^{\star}_{k}\in\mathrm{br}(\nu_{k})$ and $\nu^{\star}_{k}\in\mathrm{br}(\mu_{k})$. At this point, dividing by $K$ squaring and applying the performance difference Lemma˜K.1 that allows to decompose the global regret into a weighted some of regrets at each state, we obtain
@@
where the second last step used the Cauchy-Schwarz inequality and the last step used the Jensen’s inequality and the concavity of the square root. Analogous steps give
@@
Therefore, thanks to an application of the Azuma-Hoeffding inequality, with probability at least $1-\delta$
@@
where for the first term, we recognized that the policies updates in Algorithm˜1 can be seen as mirror descent updates (see Lemma˜E.1) and we used the standard regret bound for online mirror descent (see for example Orabona [2023]) instantiated with the following Bregman divergence $\sum_{s\in{\mathcal{S}}}D_{KL}(\mu(\cdot|s),\mu^{\prime}(\cdot|s))$ and with learning rate $\eta=\frac{2\left|{{\mathcal{S}}}\right|\log\left|{{\mathcal{A}}_{\max}}\right|}{K}$ as done in Algorithm˜1. All in all, we obtain via a union bound that with probability at least $1-4\delta$
@@
The next results shows that the policies updates used in Algorithm˜1 and Algorithm˜2 are mirror descent updates for an appropriately chosen Bregman divergence. To this end for any $p,d\in\Delta_{\mathcal{A}}$ we define the KL divergence as $KL(p,q)=\sum_{a\in\mathcal{A}}p(a)\log(p(a)/q(a))$ with the convention that $KL(p,q)=0$ if there exists an action $a$ such that $p(a)=0$ and $q(a)>0$.
Lemma E.1.
The updates used in Algorithm˜1 and Algorithm˜2, that is
@@
The Bregman divergence chosen is induced by the function $\psi(\mu)=\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\mu(a|s)\log\mu(a|s)$ sum of the negative entropy over the state space. Notice that the gradient norm tends to infinite as $\mu$ approaches the border of the policy space (i.e. some entries $\mu(a|s)$ tends to zero), that is $\lim_{\mu\rightarrow\delta\Pi}\left\|{\nabla\psi(\mu)}\right\|$. This means that the first order optimality condition implies that the derivative $F(\mu)=\left\langle{g^{\mu}_{k}},{\mu}\right\rangle+\frac{1}{\eta}\sum_{s\in{\mathcal{S}}}KL(\mu(\cdot|s),\mu_{k}(\cdot|s))$ equals zero at the minimizing policy which is $\mu_{k+1}$ by definition. Therefore, in the following we use this fact to derive the exponential weight updates used in Algorithm˜1 and 2.
@@
for some $c\in\mathbb{R}$ which ensures normalization of $\mu_{k+1}$. Therefore, inverting the last expression, we get
@@
Choosing $c\in\mathbb{R}$ to ensure that $\forall s\in\mathcal{S}$ it holds that $\sum_{a\in\mathcal{A}}\mu_{k+1}(a|s)=1$ concludes the proof. ∎
Appendix F Analysis of Algorithm˜2
This section presents the analysis of Algorithm˜2 which provides a sample complexity guarantee without requiring neither concentrability or best response oracle. See 4.2
@@ Appendix F Analysis of Algorithm˜2
The proof follows similar to the one of Theorem˜4.1 with the addition of an RL inner loop. In a first step, we first derive the same decomposition as in (5). Again, dividing by $K$ squaring, applying the performance difference, the Cauchy-Schwartz inequality leads to (6) and using Jensen’s inequality and the concavity of the square root we get
@@
where $\mu^{\star}_{k}\in\mathrm{br}(\nu_{k}).$ At this point, as we do not assume to have access to a Best Response oracle, let us introduce the sequence $\left\{{z_{k}}\right\}^{K}_{k=1}$ and $\left\{{y_{k}}\right\}^{K}_{k=1}$ produced by UCB-VI in the inner loop of Algorithm˜2. Since the stochastic reward used in the inner loop is unbiased and almost surely bounded by $2$ by Lemma˜G.7, Lemma˜G.6 run for a number of inner iterations $T=\mathcal{O}\left({\frac{\left|{{\mathcal{S}}}\right|^{3}\left|{{\mathcal{A}}_{\max}}\right|\log(1/\delta)}{(1-\gamma)^{4}\varepsilon^{2}_{\mathrm{opt}}}}\right)$ ensures that with probability $1-3\delta$
@@
Note that now we can sample $S_{k}^{\mu}\sim d^{\mu_{k},y_{k}}$ and $S_{k}^{\nu}\sim d^{z_{k},\nu_{k}}$. Therefore, we can again add and subtract the terms $\left\|{\mu_{E}(\cdot|S^{\mu}_{k})-\mu_{k}(\cdot|S^{\mu}_{k})}\right\|^{2}$ in (F) and $\left\|{\nu_{E}(\cdot|S^{\nu}_{k})-\nu_{k}(\cdot|S^{\nu}_{k})}\right\|^{2}$ in (9) we obtain that
@@
We have that (12) can be bounded analogously as in Eq.˜Martingale with $S^{\mu}_{k}\sim d^{\mu_{k},y_{k}}.$
Again, as done in the proof of Theorem˜4.1, we recognize that the policies updates performed by Algorithm˜2 are instances of online mirror descent ( see Lemma˜E.1 ). Therefore, we can bound the regret term as follows
@@
Appendix G Analysis for the RL inner loop
@@
and we prove that under the above conditions pointwise optimism hold. This point is made precise in the next Lemma.
Lemma G.2.
@@
Using the point wise optimism in Lemma G.2 and the decomposition in Lemma G.1 we have the following decomposition on the regret of UCBVI
@@
$=\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{Q_{t+1}-r+\gamma PV_{t}}\right\rangle+\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{Q_{t}-Q_{t+1}}\right\rangle$
$\leq\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{\widehat{r}_{t}+\gamma\widehat{P}_{t}V_{t}+b_{t}-r+\gamma PV_{t}}\right\rangle+\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{Q_{t}-Q_{t+1}}\right\rangle$
$\leq\frac{2}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{b_{t}}\right\rangle+\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{Q_{t}-Q_{t+1}}\right\rangle$
@@
$\frac{1}{T}\sum^{T}_{t=1}\left\langle{d^{\pi_{t}}},{Q_{t}-Q_{t+1}}\right\rangle$ $\leq\frac{1}{T}\sum^{T}_{t=1}\sum_{s,a}Q_{t}(s,a)-Q_{t+1}(s,a)$
@@
G.1 Showing validity of the bonuses
@@
For all $t\in[T]$ simultaneously, we have the following high probability upper bound
@@
where the last inequality follows with probability $1-\delta$ from an application of the Azuma Hoeffding inequality making special care of the fact that the total number of visits $N_{t}(s,a)$ is not an independent random variable with respect to the random variables of which we are computing the mean, that is $\left\{{\mathds{1}_{\left\{{S_{\tau},A_{\tau}=s,a}\right\}}}\right\}^{T}_{t=1}$. For this reason we pay the factor $\log(N_{t}(s,a)(N_{t}(s,a)+1))$ in the upper bound. We refer the reader to [Lattimore and Szepesvári, 2020, Exercise 7.1 ] for details. Therefore, by triangular inequality and a union bound over the state space.
@@
G.2 Bound the bonus sum
Lemma G.5.
@@ G.2 Bound the bonus sum
G.3 Final UCBVI bound
Lemma G.6.
Let us consider UCBVI (Algorithm 3) in an environment with a stochastic unbiased reward almost surely bounded by $2$ run for $T$ iteration with a sequence of valid bonuses $\left\{{b_{t}}\right\}^{T}_{t=1}$ specified in the statement of Lemma˜G.4, then it holds that with probability at least $1-3\delta$
@@ G.3 Final UCBVI bound
The proof follows trivially from the combination of Lemma˜G.5 and Lemma˜G.3 and a union bound over the event that the bonus are valid and the event under which the bound in Lemma˜G.5 holds . ∎
G.4 Properties of the reward estimate
Lemma G.7.
For any policy $\pi\in\Pi$ and expert policy $\pi_{E}\in\Pi$ consider a particular state $s\in{\mathcal{S}}$ and sampling $A_{E}\sim\pi_{E}(\cdot|s)$, $A^{\prime}_{E}\sim\pi_{E}(\cdot|s)$ . Then, the following facts hold true
@@ G.4 Properties of the reward estimate
$\left\|{\pi_{E}-\pi}\right\|^{2}$ $=\left\|{\pi_{E}(\cdot|s)}\right\|^{2}-2\left\langle{\pi_{E}(\cdot|s)},{x^{k}(\cdot|s)}\right\rangle+\left\|{\pi(\cdot|s)}\right\|^{2}$
$=\sum_{a}\pi_{E}(a\mid s)^{2}+\sum_{a}\pi(a\mid s)^{2}-2\sum_{a}\pi_{E}(a\mid s)\pi(a\mid s)$
$=\sum_{a}\pi_{E}(a\mid s)^{2}+\sum_{a}\pi(a\mid s)^{2}-2\mathbb{E}_{A\sim\pi_{E}(\cdot\mid s)}[\pi(a\mid s)].$
@@
$\pi_{E}^{2}(a\mid s)=\mathbb{P}(A_{E}=a)^{2}=\mathbb{P}(A_{E}=a)\mathbb{P}(A^{\prime}_{E}=a)=\mathbb{P}(A_{E}=A^{\prime}_{E})=\mathbb{E}[\mathds{1}_{\{A_{E}=A^{\prime}_{E}\}}],$
where $A_{E},A^{\prime}_{E}$ are independent samples from $\pi_{E}(\cdot\mid s).$ Therefore, we can conclude that
@@
is an unbiased estimator of $\left\|{\pi_{E}-\pi}\right\|^{2}$. The second statement is easy to show
@@
Appendix H Extension to $n$-player general-sum games
In this section, we sketch the analysis for the $n$-player general sum extension. The goal of this section is to show that the algorithm design is decentralized, meaning that it can avoid the curse of multi-agents in $n$-player general-sum Games as stated in Remark˜4.1. The idea is that the introduced algorithms keep the other players fixed, in the RL inner-loop and the BR oracle calls respectively. This results in a decentralized execution. This section starts with the introduction of $n$-player general-sum Games and all the necessary notations. Then, we show how the objective varies slightly from the one in Zero-Sum Games. Last, we give the proof sketch for the $n$-player general-sum case.
@@ Appendix H Extension to $n$-player general-sum games
$V_{i}^{\boldsymbol{\pi}}(s):=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}r_{i}(S,\boldsymbol{A})\mid S_{0}=s\right]$
$Q_{i}^{\boldsymbol{\pi}}(s,\textbf{a}):=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}r_{i}(S,\boldsymbol{A})\mid S_{0}=s,A_{0}=\textbf{a}\right].$
All other expressions are defined as in Section˜2 with the extension that the joint actions are now given by a and one fixes the policies of all players expect player $i$, i.e. $\pi_{-i}$, for the induced Games.
@@
$\leq\sum_{i=1}^{n}\left(\underbrace{V^{\pi^{\star}_{i},{\pi}_{-i}}_{i}(s_{0})-V^{\pi^{\star}_{i},{\pi}_{E_{-i}}}_{i}(s_{0})}_{\mathrm{Exploit-Gap}_{i}}+\underbrace{V^{\pi_{E_{i}},{\pi}_{E_{-i}}}_{i}(s_{0})-V^{\pi_{i},{\pi}_{-i}}_{i}(s_{0})}_{\mathrm{Value-Gap}}\right).$
The Exploit-Gap is similar to the objective analyzed in the zero-sum case with the difference that now that the policies of the other players are varying in n1n-1 cases. The Value-Gap is new and does not appear in the zero-sum case as one can again use the structure of the reward in that case. However, the latter is easy to analyze as it can be seen as a single-agent MDP with the joint policy $\boldsymbol{\pi}$ and the joint expert $\boldsymbol{\pi}_{E}$ respectively.